The equation is (2y + 1/2)^2 = 4/25. I divided the first half of the equation by 2 to get the variable alone, so my new equation is (y + 1/4)^2 = 4/25. I squared both sides, ending up with y + 1/4 = 2/5. My final answers were 13/20 and -3/20, not the answers in the book. Please help me find my error. Thank you!!
Answer by scott8148(6628)

for the sides of an equation to stay equal; whatever you do to one side, you must do to the other side

also, dividing a squared quantity by 2 is actually dividing the value by 4 (because of the squaring)

you said you squared both sides, when you actually took the square root

(2y + 1/2)^2 = 4/25

(y + 1/4)^2 = 1/25

y + 1/4 = ±1/5

y = -1/20 ___ and ___ y = -9/20